### v2

# Revisiting the Identification of the Conduct Parameter in Homogeneous Goods Markets

Abstract
We revisit the identification of the conduct parameter in homogeneous goods markets. lauIdentifying1982 argues that the conduct parameter is not identified if and only if the inverse demand function is separable, except for a specific separable function. This result has been regarded as an extension of the result in bresnahanOligopoly1982 to more general settings. However, we show that Lau’s claim is incorrect and provide a new characterization of the non-identification. Our characterization shows that a demand function with demand rotation instruments is the necessary and sufficient condition for the identification of the conduct parameter. Therefore, our result properly generalizes the role of demand rotation instruments in identifying the conduct parameter, as highlighted by bresnahanOligopoly1982, to more general settings.
Keywords: Identification, Conduct Parameter, Homogeneous Goods Market
JEL Codes: C5, C13, L1
1 Introduction
Measuring market power is an important task in several fields of economics, especially in industrial organization. While there are several measures of market power, markup is one of the most popular measures. In the last decade, the rise of markup has been a hot topic in both macroeconomics and industrial organization (see millerIndustrial2025 for the detailed survey). However, when estimating markup, most papers put restrictions on how firms compete, that is, firm conduct, and hence, the estimation result could be biased.111deloeckerRise2020 use accounting data and estimate markup based on the production function approach under the assumption of imperfect competition in an output market and perfect competition in input markets. dopperRising2021 uses Nielsen data and estimates markup under the assumption of Bertrand competition.
In the literature on empirical industrial organization, to measure markup under a flexible competition model, the conduct parameter approach has been used: porterStudy1983, genesoveTesting1998, and okazakiExcess2022 in homogeneous product markets, and millerUnderstanding2017 and cilibertoDoes2014 in differentiated product markets. In this approach, the researcher considers a generalized marginal revenue function with a conduct parameter, and estimating the conduct parameter is equivalent to estimating markup under a flexible competition model. In addition to measuring markup, recently the conduct parameter approach has received attention. weylPassThrough2013 provides a new insight between the conduct parameter and the pass-through and some empirical studies estimate the conduct parameter as an intermediate object to investigate the pass-through (millerPassthrough2017). beringeRobust2025 study the welfare analysis under imperfect competition and use the conduct parameter approach to get a robust prediction of the welfare.
A central question in this approach is the identification of the conduct parameter. As for the identification in homogeneous goods markets, there are two influential papers. First, bresnahanOligopoly1982 considers the identification of the conduct parameter and the marginal cost function in a model with linear demand and linear marginal cost. He shows that the conduct parameter and the parameters of the marginal cost function can be identified when the inverse demand function includes a demand shifter, called a demand rotation instrument. This demand rotation instrument changes the slope and intercept of the inverse demand function without changing the equilibrium point under the true model. However, the equilibrium point changes under the false model, and hence the conduct parameter can be identified by observing the equilibrium outcomes.
Second, lauIdentifying1982 considers a more general setting and investigates the conditions under which the conduct parameter is not identified. Then, he claims that the shape of the inverse demand function plays the central role in the identification by deriving the sufficient and necessary condition for non-identification:
Claim.
Under some assumptions, the conduct parameter and the marginal cost function are not identified if and only if the inverse demand function is separable in demand shifters except for a specific separable function.
The claim implies that separability is the key to the non-identification, but there is a specific separable function that always leads to the identification. The claim also implies the equivalence between non-separable demand functions and the identification. As the inverse demand function with a demand rotation instrument is not separable, this result has been regarded as an extension of bresnahanOligopoly1982’s idea in a more general setting.
The main contribution of our paper is to show that Lau’s claim is incorrect and to provide a new characterization of non-identification. Based on our characterization, we have the following result:
Result.
Under some assumptions, the conduct parameter and the marginal cost function are not identified if and only if inverse demand functions are such that demand shifters can change only the slope or the intercept of the inverse demand function, except for a specific inverse demand function.
In this result, as the demand shifters can change only the slope or the intercept of the inverse demand, the inverse demand function cannot have any demand rotation instrument that requires keeping the equilibrium quantity unchanged by changing the slope and the intercept simultaneously. Therefore, our result restricts the type of the inverse demand function for the non-identification more than the separability requirement in Lau’s claim. The contraposition of our result implies that the inverse demand function with a demand rotation instrument is the necessary and sufficient condition for the identification, and hence our result clearly reinforces the role of demand rotation instruments in the identification of the conduct parameter and indicates that the idea in bresnahanOligopoly1982 is crucial for the identification of the conduct parameter even in more general settings.
Why is Lau’s result incorrect? The idea of Lau’s proof is the following: let $\mathcal{E}=(\theta,MC)$ be the pair of the conduct parameter and the marginal cost function, and let $\mathcal{E}^{*}=(\theta^{*},MC^{*})$ be an alternative pair. The non-identification implies that given demand shifters and cost shifters, both $\mathcal{E}$ and $\mathcal{E}^{*}$ lead to the same equilibrium quantity. Lau shows that the non-identification implies that there is a transformation for the equilibrium conditions between $\mathcal{E}$ and $\mathcal{E}^{*}$, that is, an equilibrium quantity satisfied in $\mathcal{E}$ is also satisfied in $\mathcal{E}^{*}$, which implies that observational equivalence holds. However, we point out that Lau’s transformation does not work properly. When a demand rotation instrument varies so as not to change the equilibrium in $\mathcal{E}$, to keep observational equivalence, his transformation requires that the marginal cost function $MC^{*}$ should depend on the demand shifter, which violates the exclusive demand shifters assumption. Therefore, Lau’s transformation does not work properly.
To resolve the problem in Lau’s transformation, we construct a new transformation of the marginal cost function between $MC$ and $MC^{*}$ based on the definition of the non-identification. Under our transformation, the transformed marginal cost function still could depend on demand shifters. Thus, to make the transformation valid, we restrict the functional form of the inverse demand function such that the transformation is independent of the demand shifter. The restriction leads to a differential equation of the inverse demand function, and the solution of the differential equation leads to the inverse demand function without demand rotation instrument.
Another contribution of our paper is that Lau’s claim assumes that the inverse demand function is already identified. However, we show that the inverse demand function that leads to the non-identification of the conduct parameter and the marginal cost function is not identified. In other words, when the observed data is generated by the inverse demand function that leads to the non-identification of the conduct parameter and the marginal cost function, we cannot identify the inverse demand function, and hence, all model components of the conduct parameter model cannot be identified. In contrast, we show that demand rotation instruments do not help to identify the inverse demand function. Therefore, we conclude that when an inverse demand function with demand rotation instruments is identified, the conduct parameter and the marginal cost function can be identified.
The rest of the paper is organized as follows. Section 2 describes the setting. Section 3 discusses Lau’s result and our main result. Subsection 3.3 and Subsection 3.4 provide a new characterization of non-identification and show the necessary and sufficient condition for non-identification. Section 4 considers whether the inverse demand function that leads to non-identification can be identified. Section 5 discusses some criticisms of the conduct parameter approach. Section 6 concludes. Appendix includes the omitted proofs in the main text and a summary of goldmanNote1964.
2 Setting
2.1 Conduct Parameter Model
Consider a homogeneous product market with an aggregate inverse demand and an aggregate marginal cost function denoted as $P(Q,\bm{X}^{d})$ and $MC(Q,\bm{X}^{s})$, respectively, where $Q$ is the aggregate product quantity, $\bm{X}^{d}$ and $\bm{X}^{s}$ are the vector of demand shifters and cost shifters, respectively. Note that every vector is represented by a bold letter, and its element or scalar value is represented by a non-bold letter. Let $K_{d}$ and $K_{s}$ be the dimension of $\bm{X}^{d}$ and $\bm{X}^{s}$, respectively.
Given the demand shifter $\bm{X}^{d}$ and the cost shifter $\bm{X}^{s}$, the equilibrium quantity $Q^{e}$ is defined as the solution to the following equation:
$P(Q,\bm{X}^{d})+\theta Q\frac{\partial P}{\partial Q}(Q,\bm{X}^{d})=MC(Q,\bm{X}^{s}),$ (1)
where $\theta\in[0,1]$ is called the conduct parameter. By rewriting the equilibrium condition (1), we have
$\theta=\frac{P-MC}{P}\varepsilon^{d},$ (2)
where $\varepsilon^{d}$ is the price elasticity of demand. Therefore, the conduct parameter is also regarded as the elasticity-adjusted Lerner index and represents the degree of market power of the firms. Depending on the value of $\theta$, the condition can represent the first-order condition of several models between perfect competition ($\theta=0$) and joint-profit maximization ($\theta=1$). [^2] It can also nest Cournot competition when $\theta=1/N$ under some marginal cost function such as constant marginal cost and linear marginal cost. In general, this holds when the aggregation of each firm’s first-order condition results in the aggregate first-order condition. Therefore, the left-hand side is regarded as a generalized marginal revenue with $\theta$, and hence the condition is a generalized first-order condition with the conduct parameter.
For the identification analysis, we put some restrictions. First, we assume that the demand shifter and the cost shifter are mutually exclusive:
Assumption 1.
The demand shifter $\bm{X}^{d}$ and the cost shifter $\bm{X}^{s}$ are exclusive; that is, there is no common variable in $\bm{X}^{d}$ and $\bm{X}^{s}$, and $\bm{X}^{d}$ affects the inverse demand function but not the marginal cost function, and $\bm{X}^{s}$ affects the marginal cost function but not the inverse demand function.
We can include some common variables in $\bm{X}^{d}$ and $\bm{X}^{s}$, but in that case, we must have at least one exclusive demand shifter and one exclusive cost shifter. Hereafter, we assume that there is no common variable in $\bm{X}^{d}$ and $\bm{X}^{s}$, for simplicity. The next assumption restricts the effectiveness of the shifters:
Assumption 2.
The given demand shifters and the cost shifters should affect the inverse demand function and the marginal cost function. Formally, we should have that for all $i=1,\ldots,K_{d}$ and $j=1,\ldots,K_{s}$,
$\frac{\partial P}{\partial X^{d}_{i}}(Q,\bm{X}^{d})\neq 0\quad\text{and}\quad\frac{\partial MC}{\partial X^{s}_{j}}(Q,\bm{X}^{s})\neq 0$ (3)
for any $Q$, $\bm{X}^{d}$, and $\bm{X}^{s}$.
When this condition is not met, there could be areas where the inverse demand function and the marginal cost function cannot be identified. For example, when a demand shifter does not change the inverse demand function for some interval of the demand shifter, the equilibrium quantity is not affected by the change in the demand shifter. Then, there is no variation to identify the inverse demand function on the interval. Because our identification result for the conduct parameter and the marginal cost function presumes the identification of the inverse demand function. Thus, to guarantee the global identification of the inverse demand function, the assumption is necessary.
Next, we impose the following assumption on the differentiability of the inverse demand function and the marginal cost function:
Assumption 3.
The inverse demand function is three times continuously differentiable, and the marginal cost function is twice continuously differentiable.
As we will see, lauIdentifying1982 assumes twice-continuous differentiability of the inverse demand function and the marginal cost function. Hence our result requires a stronger assumption on the inverse demand function. However, we will justify our assumption by showing that Lau’s claim also needs three-times continuous differentiability.
Finally, we impose the equilibrium existence condition:
Assumption 4.
Given an inverse demand function, a conduct parameter, and a marginal cost function, and given $\bm{X}^{d}$ and $\bm{X}^{s}$, the derivative of the equilibrium condition with respect to $Q$ is not zero, that is,
$(1+\theta)\frac{\partial P}{\partial Q}(Q,\bm{X}^{d})+\theta Q\frac{\partial^{2}P}{\partial Q^{2}}(Q,\bm{X}^{d})-\frac{\partial MC}{\partial Q}(Q,\bm{X}^{s})$ $\neq 0.$ (4)
Note that as the above equation consists of the derivative of the inverse demand and the marginal cost, by Assumption 3, the derivative of the equilibrium condition with respect to $Q$ is finite, and hence the left-hand side is a finite value for any $Q$, $\bm{X}^{d}$, and $\bm{X}^{s}$.
2.2 The Data Generation Process
Suppose that the researcher observes the aggregate price $P$ and the aggregate quantity $Q$, and the vector of exogenous variables $\bm{X}^{d}$ and $\bm{X}^{s}$. Assume that the data is generated through the equilibrium condition (1). We assume that from the data, the reduced form of the equilibrium quantity and the equilibrium price are identified:
Assumption 5.
The reduced form of the equilibrium quantity and the equilibrium price, defined as
$Q^{e}=h_{q}(\bm{X}^{d},\bm{X}^{s}),\quad P^{e}=h_{p}(\bm{X}^{d},\bm{X}^{s}),$ (5)
are identified.
The identification of the reduced-forms can be done simply due to the variation in the demand shifters and the cost shifters.
2.3 Definitions of Non-identification and Identification
Now, we introduce the definition of the identification problem in this setting. lauIdentifying1982 considers the identification of the conduct parameter and the marginal cost function. While our interest is the identification, we take an indirect approach and specify the conditions under which the model is not identified. The definition of the non-identification is as follows:
Definition 1.
The conduct parameter and the marginal cost function are said to be non-identified if there are two distinct pairs of conduct parameter and marginal cost function, denoted as $(\theta,MC)$ and $(\theta^{*},MC^{*})$ such that the corresponding reduced-form equilibrium quantity are identical for demand shifter and cost shifter: $h_{q}(\bm{X}^{d},\bm{X}^{s})=h_{q}^{*}(\bm{X}^{d},\bm{X}^{s})$ for any $\bm{X}^{d}$ and $\bm{X}^{s}$. In other words, for each $\bm{X}^{d}$ and $\bm{X}^{s}$, there is an equilibrium quantity $Q^{e}$ that satisfies both
$P(Q^{e},\bm{X}^{d})+\theta Q^{e}\frac{\partial P}{\partial Q}(Q^{e},\bm{X}^{d})$ $=MC(Q^{e},\bm{X}^{s}).$ (6)
and
$P(Q^{e},\bm{X}^{d})+\theta^{*}Q^{e}\frac{\partial P}{\partial Q}(Q^{e},\bm{X}^{d})$ $=MC^{*}(Q^{e},\bm{X}^{s}).$ (7)
Then, by taking the contraposition of Definition 1, we can characterize the identification of the conduct parameter and the marginal cost function:
Corollary 1.
The conduct parameter and the marginal cost are said to be identified if for any two distinct pairs of a conduct parameter and a marginal cost function, there exists a pair of demand shifters $\tilde{\bm{X}}^{d}$ and $\tilde{\bm{X}}^{s}$ where the equilibrium quantity $Q^{e}$ is different, that is, $h_{q}(\tilde{\bm{X}}^{d},\tilde{\bm{X}}^{s})\neq h_{q}^{*}(\tilde{\bm{X}}^{d},\tilde{\bm{X}}^{s})$.
Note that the non-identification and the identification assume the identification of the inverse demand function. In other words, we consider the identification of the conduct parameter and the marginal cost function given an identified inverse demand function. While it should be discussed whether the inverse demand function can be identified from the data, we will discuss the identification of the inverse demand function later.
3 Main Result: Characterization of the Non-identification
3.1 Lau’s Claim
Based on the setting in Section 2, lauIdentifying1982 shows the equivalence between the separability of the inverse demand function and the non-identification of the conduct parameter and the marginal cost function. The following claim is the quote of Theorem 1 in lauIdentifying1982.
Claim 1.
Under the assumption that the inverse demand function and the marginal cost functions are twice continuously differentiable, the index of competitiveness $\theta$ cannot be identified from data on industry price and quantity and other exogenous variables alone if and only if the industry inverse demand function is separable in the demand shifter, that is,
$P(Q,\bm{X}^{d})=P(Q,r(\bm{X}^{d})),$ (8)
but does not take the form
$P(Q,r(\bm{X}^{d}))=Q^{-\frac{1}{\theta}}r(\bm{X}^{d})+s(Q).$ (9)
The result says that the shape of the inverse demand function is important for the identification. When the inverse demand function is separable in demand shifters, the conduct parameter is not identified. An inverse demand function satisfying (8) has weak separability defined in goldmanNote1964.  [^3] See Definition B.1 in Appendix B. Note that the separability in goldmanNote1964 is defined only when the dimension of the demand shifter is greater than two. When the demand shifter is a scalar, we cannot apply the definition of separability in goldmanNote1964.
To understand the intuition of Lau’s claim, let us consider an example in bresnahanOligopoly1982. In his paper, he considers a market with a linear inverse demand and a linear marginal cost. To break the separability, he considers a linear inverse demand function such that
$P(Q,\bm{X}^{d})=\alpha_{0}-\alpha_{1}Q+\alpha_{2}X^{d}_{1}+\alpha_{3}QX^{d}_{1}+\alpha_{4}X^{d}_{2},$ (10)
where $X^{d}_{1}$ is called a demand rotation instrument because it can change the slope and the intercept of the inverse demand function without changing the equilibrium quantity. While Bresnahan shows the identification under the linear inverse demand function, Lau’s claim extends Bresnahan’s result to more general settings.
The claim also emphasizes that the dimension of the demand shifter is important for the identification. When the demand shifter is a scalar, that is, when the dimension of the demand shifter is one, the inverse demand function (8) nests any inverse demand with a scalar demand shifter because under a scalar demand shifter $X^{d}$, the change of the demand shifter through $r(X^{d})$ is regarded as the same as the change in $X^{d}$. Thus, the conduct parameter can be identified only when the inverse demand function takes the form (9) when the demand shifter is a scalar.
While Lau’s claim has been regarded as the extension of bresnahanOligopoly1982 to more general settings, an important observation is that introducing demand rotation instruments breaks the separability, but the separability does not exclude demand rotation instruments. For example, consider an inverse demand function given as
$P(Q,r(\bm{X}^{d}))=-Qr(\bm{X}^{d})+Q+r(\bm{X}^{d})=r(\bm{X}^{d})(1-Q)-Q,$ (11)
where the range of $r$ is greater than one.
First, based on the definition of weak separability in goldmanNote1964, the function is verified to be separable because for any $i$ and $j$:
$\frac{\partial}{\partial Q}\left(\frac{\frac{\partial P}{\partial X^{d}_{i}}(Q,\bm{X}^{d})}{\frac{\partial P}{\partial X^{d}_{j}}(Q,\bm{X}^{d})}\right)=\frac{\partial}{\partial Q}\left(\frac{r_{i}(\bm{X}^{d})(1-Q)}{r_{j}(\bm{X}^{d})(1-Q)}\right)=\frac{\partial}{\partial Q}\frac{r_{i}(\bm{X}^{d})}{r_{j}(\bm{X}^{d})}=0,$ (12)
where $r_{i}(\bm{X}^{d})\equiv\frac{\partial r(\bm{X}^{d})}{\partial\bm{X}^{d}_{i}}$.
Second, we can verify that the demand shifters can work as demand rotation instruments. Given the inverse demand function, the marginal revenue under $\theta$ is written as
$MR(Q,X^{d};\theta)=-r(\bm{X}^{d})(1-(1+\theta)Q)+Q(1+\theta).$ (13)
When the quantity is $Q^{\prime}=\frac{1}{1+\theta}$, the marginal revenue is not affected by the change in the demand shifter $X^{d}$ because it is equal to one. Suppose that there is a marginal cost function $MC$ where $Q^{\prime}$ holds as an equilibrium. Then, as the marginal revenue is a constant in the demand shifter, any change in the demand shifter does not affect the equilibrium quantity. In contrast, suppose that under another conduct parameter $\theta^{*}$ and another marginal cost function $MC^{*}$, $Q^{\prime}$ holds as an equilibrium, that is, $Q^{\prime}$ satisfies
$r(\bm{X}^{d})(1-(1+\theta^{*})Q^{\prime})+Q^{\prime}(1+\theta^{*})=MC^{*}(Q^{\prime},\bm{X}^{s}).$ (14)
Here, the marginal revenue under $\theta^{*}$ is a function of the demand shifter at $Q^{\prime}$. Thus, the change in $\bm{X}^{d}$ leads to a different equilibrium quantity from $Q^{\prime}$, which implies the violation of the observational equivalence. Therefore, the demand shifter $\bm{X}^{d}$ can work as the demand rotation instrument at $Q^{\prime}$ because the demand shifter changes the slope and the intercept of the inverse demand function simultaneously without changing the equilibrium quantity under the true conduct but the change leads to a different equilibrium quantity under false conduct.
While the example admits a separable inverse demand function with demand rotation instrument, unfortunately, Lau does not clearly explain how this situation is treated in his claim, and hence, it is not clear why not demand rotation instruments but the separability of the inverse demand function is the key to identification.
3.2 Main Result
We are going to show the following theorem, which is the formal version of the result in the introduction:
Theorem 1.
Given Assumption 1, 2, 3, and 4, the conduct parameter $\theta$ cannot be identified from data on price, quantity, and other exogenous variables alone if and only if the industry inverse demand function is given as
$P(Q,\bm{X}^{d})=Q^{\alpha}r(\bm{X}^{d})+s(Q)$ (15)
where $\alpha\neq-\frac{1}{\theta}$ is a constant, and $r:\mathbb{R}^{K^{d}}\rightarrow\mathbb{R}$ and $s:\mathbb{R}_{+}\rightarrow\mathbb{R}$ are twice continuously differentiable functions
As in Lau’s claim, the demand shifters affect the inverse demand function only through a function $r$. Therefore, when the demand shifter is a vector, the inverse demand function is separable. However, our result tells more about the non-identification holds under what type of separable function and has a clear economic interpretation. Again, our result has a special case where the identification always holds ($\alpha=-\frac{1}{\theta}$). As we will discuss, the identification is always achieved because any demand shifter works as a demand rotation instrument in this case. Except in this case, observe that the demand shifter changes only the slope of the inverse demand function when $\alpha\neq 0$ or changes only the intercept of the inverse demand function when $\alpha=0$. Recall that a demand rotation instrument can change the intercept and the slope of the inverse demand function simultaneously without changing the equilibrium quantity. Therefore, we cannot have any demand rotation instrument under (15). The result implies that the identification holds if and only if the inverse demand function has demand shifters that work as demand rotation instruments. This clearly reinforces the idea of bresnahanOligopoly1982 using demand rotation instruments to identify the conduct parameter in more general settings. It also reinforces the sufficiency of demand rotation instruments for the identification in the model with a linear inverse demand and a linear marginal cost as in bresnahanOligopoly1982 and matsumuraResolving2023, extends it to more general settings and shows that the condition is not only sufficient but also necessary.
Hereafter, we provide the proof of Theorem 1. The proof of Theorem 1 has the following steps. First, we derive the necessary condition for non-identification based on Definition 1. By rewriting Definition 1, we obtain the transformation that transforms the derivatives of $MC$ into the derivatives of another marginal cost function $MC^{*}$ that leads to observational equivalence given an inverse demand function (Lemma 2). However, the transformation is not valid because it consists of the derivatives of the inverse demand function, that is, it depends on the demand shifter $\bm{X}^{d}$. Then, without any treatment for the demand shifters, $MC^{*}$ depends on the demand shifter $\bm{X}^{d}$, which violates Assumption 1. Therefore, to make the transformation valid, we need to remove the effect of the demand shifter from the transformation, which puts restrictions on the inverse demand function. Then, the restriction leads to a differential equation that characterizes the inverse demand function that leads to non-identification (Proposition 1). Then, to show sufficiency, under the inverse demand function characterized in Proposition 1, we construct a transformation that leads to observational equivalence (Proposition 2).
3.3 Necessary Condition for the Non-identification
First, we characterize the non-identification condition based on Definition 1. The characterization is based on the fact that observational equivalence implies that the reduced forms of the equilibrium quantity $h_{q}$ and $h_{q}^{*}$ are identical for any $\bm{X}^{d}$ and $\bm{X}^{s}$. Therefore, its derivative with respect to the demand and cost shifters, $\nabla h_{q}$ and $\nabla h_{q}^{*}$, should also be identical. Then, by applying the implicit function theorem to the equilibrium condition, we can compute $\nabla h_{q}$ and $\nabla h_{q}^{*}$. The following lemma characterizes the non-identification condition:
Lemma 1.
Non-identification implies that for any $\bm{X}^{d}$, $\bm{X}^{s}$, and $Q^{e}$ under these exogenous variables, we have for $i=1,\ldots,K_{d}$,
$\frac{\partial P}{\partial X^{d}_{i}}(Q^{e},\bm{X}^{d})+\theta^{*}Q^{e}\frac{\partial^{2}P}{\partial X^{d}_{i}\partial Q}(Q^{e},\bm{X}^{d})$ (16)
$=\lambda(Q^{e},\bm{X}^{d},\bm{X}^{s})\left[\frac{\partial P}{\partial X^{d}_{i}}(Q^{e},\bm{X}^{d})+\theta Q^{e}\frac{\partial^{2}P}{\partial X^{d}_{i}\partial Q}(Q^{e},\bm{X}^{d})\right],$ (17)
and for $j=1,\ldots,K_{s}$,
$\frac{\partial MC^{*}}{\partial X^{s}_{j}}(Q^{e},\bm{X}^{s})=\lambda(Q^{e},\bm{X}^{d},\bm{X}^{s})\frac{\partial MC}{\partial X^{s}_{j}}(Q^{e},\bm{X}^{s}),$ (18)
where $\lambda(Q^{e},\bm{X}^{d},\bm{X}^{s})$ is defined as
$\lambda(Q^{e},\bm{X}^{d},\bm{X}^{s})\equiv\frac{(1+\theta^{*})\frac{\partial P}{\partial Q}(Q^{e},\bm{X}^{d})+\theta^{*}Q^{e}\frac{\partial^{2}P}{\partial Q^{2}}(Q^{e},\bm{X}^{d})-\frac{\partial MC^{*}}{\partial Q}(Q^{e},\bm{X}^{s})}{(1+\theta)\frac{\partial P}{\partial Q}(Q^{e},\bm{X}^{d})+\theta Q^{e}\frac{\partial^{2}P}{\partial Q^{2}}(Q^{e},\bm{X}^{d})-\frac{\partial MC}{\partial Q}(Q^{e},\bm{X}^{s})}.$ (19)
See Appendix A for the detailed proof. Equation (17) and (18) imply that when non-identification holds, the derivative of the marginal revenue with $\theta$ can be transformed into the derivative of the marginal revenue with $\theta^{*}$, and the derivative of the marginal cost $MC$ with respect to $\bm{X}^{s}$ can be transformed into the derivative of the marginal cost $MC^{*}$ with respect to $\bm{X}^{s}$ by $\lambda(Q^{e},\bm{X}^{d},\bm{X}^{s})$.
These transformations cannot always be valid because the transformed marginal revenue is affected by the cost shifter $\bm{X}^{s}$ and the transformed marginal cost is affected by the demand shifter $\bm{X}^{d}$ through $\lambda$. These dependencies are not allowed by Assumption 1. Thus, to consider a valid transformation, we further rewrite (17) and (18). The next lemma provides a transformation of the derivative of $MC$ and $MC^{*}$ with respect to $Q$ and $\bm{X}^{s}$ that leads to observational equivalence:
Lemma 2.
For any $Q^{e}$, $\bm{X}^{d}$, and $\bm{X}^{s}$, non-identification implies that the derivative of marginal cost $MC$ can be transformed into the derivative of marginal cost $MC^{*}$: for $i=1,\ldots,K_{d}$, and $j=1,\ldots,K_{s}$,
$\frac{\partial MC^{*}}{\partial Q}(Q^{e},\bm{X}^{s})=D_{i}(Q^{e},\bm{X}^{d})+C_{i}(Q^{e},\bm{X}^{d})\frac{\partial MC}{\partial Q}(Q^{e},\bm{X}^{s}),$ (20)
and
$\frac{\partial MC^{*}}{\partial X^{s}_{j}}(Q^{e},\bm{X}^{s})=C_{i}(Q^{e},\bm{X}^{d})\frac{\partial MC}{\partial X^{s}_{j}}(Q^{e},\bm{X}^{s}).$ (21)
where $C_{i}(Q,\bm{X}^{d})$ and $D_{i}(Q,\bm{X}^{d})$ are defined as
$C_{i}(Q,\bm{X}^{d})\equiv\frac{\frac{\partial P}{\partial X^{d}_{i}}(Q,\bm{X}^{d})+\theta^{*}Q\frac{\partial^{2}P}{\partial X^{d}_{i}\partial Q}(Q,\bm{X}^{d})}{\frac{\partial P}{\partial X^{d}_{i}}(Q,\bm{X}^{d})+\theta Q\frac{\partial^{2}P}{\partial X^{d}_{i}\partial Q}(Q,\bm{X}^{d})},$ (22)
and
$D_{i}(Q,\bm{X}^{d})$ $\equiv\frac{\theta^{*}-\theta}{\frac{\partial P}{\partial X^{d}_{i}}(Q,\bm{X}^{d})+\theta Q\frac{\partial^{2}P}{\partial X^{d}_{i}\partial Q}(Q,\bm{X}^{d})}\Bigg[\frac{\partial P}{\partial Q}(Q,\bm{X}^{d})\frac{\partial P}{\partial X^{d}_{i}}(Q,\bm{X}^{d})$ (23)
$\quad+Q\frac{\partial^{2}P}{\partial Q^{2}}(Q,\bm{X}^{d})\frac{\partial P}{\partial X^{d}_{i}}(Q,\bm{X}^{d})-Q\frac{\partial P}{\partial Q}(Q,\bm{X}^{d})\frac{\partial^{2}P}{\partial X^{d}_{i}\partial Q}(Q,\bm{X}^{d})\Bigg]$ (24)
See Appendix A for the detailed proof. The subscript $i$ in $C_{i}$ and $D_{i}$ indicates that $C_{i}$ and $D_{i}$ consist of the derivative of the inverse demand function with respect to $X^{d}_{i}$. Unlike Lemma 1, we have only the transformation relating to marginal cost, and hence we can see the potential obstacle in this transformation more clearly: for the transformations to hold universally for any $Q^{e}$ and $\bm{X}^{s}$, they must be independent of $\bm{X}^{d}$. If the dependence holds, while a demand rotation instrument does not change the equilibrium quantity $Q^{e}$, it changes $MC^{*}$ via $C_{i}$ and $D_{i}$. However, this means that the marginal cost funciton is af funciton of the demand shifter $\bm{X}^{d}$, which is prohibited by Assumption 1. Therefore, in order for the transformations to be valid, both $C_{i}$ and $D_{i}$ must be independent of $\bm{X}^{d}$.
To see when the terms $C_{i}$ and $D_{i}$ independent of $\bm{X}^{d}$, we take the derivative of $C_{i}$ and $D_{i}$ with respect to $X^{d}_{j}$ and put them as zero for all $j=1,\ldots,K^{d}$. A concern is that, as $C_{i}$ and $D_{i}$ already consist of the second-order derivative of the inverse demand function, taking the derivative of $C_{i}$ and $D_{i}$ leads to three-times derivative of the inverse demand function. This is why we need a stronger restriction (Assumption 3) on the inverse demand function than Lau considers (the twice-continuous differentiability assumption in Lau’s claim).
Case: The denominators of $C_{i}$ and $D_{i}$ are zero for some demand shifters
Before we characterize the inverse demand function that makes $C_{i}$ and $D_{i}$ independent of $\bm{X}^{d}$, we first check when $C_{i}$ and $D_{i}$ are not well-defined. This implies that (22) and (24) are violated, and hence the identification holds by Corollary 1. All derivatives in $C_{i}$ and $D_{i}$ are finite by Assumption 3, and hence the numerator and the denominator of $C_{i}$ and $D_{i}$ cannot take a infinite value. In this case, no matter what value of the numerator of $C_{i}$ and $D_{i}$ is, (22) and (24) are violated when at least one of the denominators of $C_{i}$ and $D_{i}$ for $i=1,\ldots,K_{d}$ is zero. In this case, $C_{i}$ and $D_{i}$ will be infinite or indeterminate.
When the denominator of $C_{i}$ or $D_{i}$ is zero for some $i$, we have that for some $(\tilde{Q},\tilde{\bm{X}}^{d})$,
$\frac{\partial P}{\partial X^{d}_{i}}(\tilde{Q},\tilde{\bm{X}}^{d})+\theta\tilde{Q}\frac{\partial^{2}P}{\partial X^{d}_{i}\partial Q}(\tilde{Q},\tilde{\bm{X}}^{d})=0.$ (25)
Let $\mathcal{I}$ be the set of indices of the demand shifters where (25) holds. Note that (25) can be rewritten as
$\frac{\partial}{\partial X^{d}_{i}}\left(P(\tilde{Q},\tilde{\bm{X}}^{d})+\theta\tilde{Q}\frac{\partial P}{\partial Q}(\tilde{Q},\tilde{\bm{X}}^{d})\right)=0,$ (26)
which is the derivative of the marginal revenue under $\theta$ with respect to $X^{d}_{i}$. Therefore, this implies that the marginal revenue under $\theta$ is not affected by the change in $X^{d}_{i}$ at $(\tilde{Q},\tilde{\bm{X}}^{d})$. This also implies that the equilibrium quantity is not affected by the change in $\tilde{X}^{d}_{i}$ at $(\tilde{Q},\tilde{\bm{X}}^{d})$. Thus, $\tilde{X}^{d}_{i}$ has the same role as a demand rotation instrument at $(\tilde{Q},\tilde{\bm{X}}^{d})$ under $\theta$.
The non-identification implies that (25) does not hold for any $Q^{e}$ and $\bm{X}^{d}$. Once we have $(\tilde{Q},\tilde{\bm{X}}^{d})$ for some $i$ where (25) holds, it is enough for the identification. While it is hard to characterize the inverse demand function that satisfies (25) for some $(\tilde{Q},\tilde{\bm{X}}^{d})$ and for some $i$, we can characterize the inverse demand function that satisfies (25) for all $(Q,\bm{X}^{d})$. Note that (25) is a partial differential equation, and we can be solved analytically. Thus, we can characterize the inverse demand function that always leads to identification for any $Q$ and $\bm{X}^{d}$:
Lemma 3.
Suppose that $\theta\neq 0$. Then, the conduct parameter and the marginal cost function are identified when the inverse demand function is such that
$P(Q,\bm{X}^{d})=Q^{-\frac{1}{\theta}}r(\bm{X}^{d})+s(Q,\bm{X}^{d}_{-\mathcal{I}}),$ (27)
where $\bm{X}^{d}_{-\mathcal{I}}$ is the vector of demand shifters whose index is not in $\mathcal{I}$, and $r:\mathbb{R}^{K_{d}}\rightarrow\mathbb{R}$ and $s:\mathbb{R}^{|\bm{X}^{d}_{-\mathcal{I}}|+1}$ are twice continuously differentiable functions.
See Appendix A for the proof. Under (27), the equilibrium condition (6) is given as
$s(Q,\bm{X}^{d}_{-\mathcal{I}})+\theta Qs^{\prime}(Q,\bm{X}^{d}_{-\mathcal{I}})=MC(Q,\bm{X}^{s}),$ (28)
Therefore, both imply that when (25) holds, the equilibrium quantity is not affected by the change in $X^{d}_{i}$, and hence $X^{d}_{i}$ always works as a demand rotation instrument under $\theta$. On the other hand, for any other $\theta^{*}\neq\theta$, the equilibrium condition becomes
$Q^{-\frac{1}{\theta}}r(\bm{X}^{d})\left(1-\frac{\theta}{\theta^{*}}\right)+s(Q,\bm{X}^{d}_{-\mathcal{I}})+Qs^{\prime}(Q,\bm{X}^{d}_{-\mathcal{I}})=MC(Q,\bm{X}^{s}).$ (29)
As $r(\bm{X}^{d})$ is a function of $X^{d}_{i}$ by Assumption 1, the equilibrium condition implies that the equilibrium quantity should depend on $X^{d}_{i}$ under $\theta^{*}$. Therefore, while the change in $X^{d}_{i}$ does not change the equilibrium quantity under $\theta$, it changes under $\theta^{*}$, which violates observational equivalence. Therefore, the conduct parameter and the marginal cost function are always identified.
Recall that Lau’s claim includes an edge case where a separable inverse demand function leads to the identification. The above lemma also includes his edge case when $\mathcal{I}=\emptyset$, that is, all demand shifters work as a demand rotation instrument. We will see that (27) is also the edge case in our result.
Case: The denominators of $C_{i}$ and $D_{i}$ are not zero
Now, suppose that (25) does not hold for any $Q$ and $\bm{X}^{d}$ and for all demand shifters. Then, we take the derivative of $C_{i}$ and $D_{i}$ with respect to $X^{d}_{j}$ for all $j=1,\ldots,K_{d}$ and put the derivative being equal to zero. This process leads to differential equations of the inverse demand functions. Fortunately, these differential equations have analytical solutions, and we can characterize an inverse demand function that satisfies the independence of $C_{i}$ and $D_{i}$ of the demand shifter:
Proposition 1.
The non-identification of the conduct parameter and the marginal cost function implies (25) does not hold for all $i=1,\ldots,K_{d}$, and $C_{i}(Q,\bm{X}^{d})$ and $D_{i}(Q,\bm{X}^{d})$ are independent of $\bm{X}^{d}$ for all $i$. The independence implies that the inverse demand function is given as
$P(Q,\bm{X}^{d})=Q^{\alpha}r(\bm{X}^{d})+s(Q)$ (30)
where $\alpha\neq-\frac{1}{\theta}$ is a constant, and $r:\mathbb{R}\rightarrow\mathbb{R}$ and $s:\mathbb{R}_{+}\rightarrow\mathbb{R}$ are twice continuously differentiable functions.
See Appendix A for the proof. This proposition characterizes the necessary condition of the non-identification. As we have discussed in the introduction, this inverse demand function does not allow for any demand rotation instrument. Recall that bresnahanOligopoly1982 provides the idea of using a demand rotation instrument to identify the conduct parameter, matsumuraResolving2023 formalize his idea by deriving a sufficient condition that guarantees the inverse demand function includes demand rotation instruments for the non-identification in Bresnahan’s setting. The contraposition of their sufficient condition implies that the non-identification holds only if the inverse demand function does not include any demand shifter. As the current environment is more general than bresnahanOligopoly1982 and matsumuraResolving2023, Proposition 2 is the refinement of the result in matsumuraResolving2023 in general settings.
A difference from Lau’s claim is that the conduct parameter can be identified even when the demand shifter is a scalar as long as the inverse demand function includes a demand rotation instrument. For example, consider an inverse demand function with a scalar demand rotation instrument:
$P(Q,X^{d}_{1})=\alpha_{0}+\alpha_{1}Q+\alpha_{2}X^{d}_{1}+\alpha_{3}QX^{d}_{1}.$ (31)
This case does not correspond to (15). Note also that $C_{1}$ and $D_{1}$ are given as
$C_{1}(Q,X^{d}_{1})=\frac{\alpha_{2}+\theta\alpha_{3}Q}{\alpha_{2}+\theta^{*}\alpha_{3}Q}\quad\text{ and }\quad D_{1}(Q,X^{d}_{1})=(\theta^{*}-\theta)(-\alpha_{1}+\alpha_{3}X^{d}_{1}).$ (32)
This implies that the transformation (20) is not valid for any marginal cost function because $D_{1}(Q,X^{d}_{1})$ depends on $X^{d}_{1}$. Therefore, any $MC^{*}$ that leads to observational equivalence with $MC$ should depend on the demand shifter, under which the identification is possible.
3.4 Sufficient Condition for Non-identification
We turn to check the sufficiency of the inverse demand function characterized in Proposition 1. To see whether the inverse demand function (15) can lead to the non-identification, we check whether there exists a transformation between the equilibrium conditions under $\mathcal{E}=(\theta,MC)$ and $\mathcal{E}^{*}=(\theta^{*},MC^{*})$. In other words, we check whether there exists a transformation $T$ such that
$MC^{*}(Q,\bm{X}^{s})=T\left(MC(Q,\bm{X}^{s})\right).$ (33)
and
$P(Q,\bm{X}^{d})+\theta^{*}Q\frac{\partial P}{\partial Q}(Q,\bm{X}^{d})=T\left(P(Q,\bm{X}^{d})+\theta Q\frac{\partial P}{\partial Q}(Q,\bm{X}^{d})\right).$ (34)
Suppose that there exists such a transformation $T$ and that an equilibrium quantity $Q^{e}$ satisfies the equilibrium condition under $\mathcal{E}$. Under $\mathcal{E}^{*}$, we have that
$P(Q^{e},\bm{X}^{d})+\theta^{*}Q^{e}\frac{\partial P}{\partial Q}(Q^{e},\bm{X}^{d})-MC^{*}(Q^{e},\bm{X}^{s})$ (35)
[…]
The inside of $T$ is the equilibrium condition under $\mathcal{E}$. Therefore, when $T(0)=0$, the above implies that $Q^{e}$ also satisfies the equilibrium condition under $\mathcal{E}^{*}$, and hence the non-identification holds.
In fact, under the inverse demand function (15), we can construct the transformation $T$ by using (20). Therefore, we can have the following sufficient condition for the non-identification:
Proposition 2.
Given the inverse demand function (15) where $\alpha\neq-\frac{1}{\theta}$, there exists a transformation $T$ between the equilibrium conditions under $\mathcal{E}$ and $\mathcal{E}^{*}$. Therefore, non-identification of the conduct parameter and the marginal cost function holds.
The proof is given in Appendix A. While Lau shows the separable inverse demand function is also a sufficient condition of the non-identification, Lau’s sufficiency proof is incomplete because it can be verified that his proof implicitly assumes that there exists a transformation $T$ under a separable inverse demand function (8). However, the separability itself does not guarantee the existence of the transformation. In contrast, Proposition 2 shows that the inverse demand function without any demand rotation instrument can lead to a transformation that leads to non-identification.
3.5 Discussion
Rethinking the role of demand rotation instrument:
Proposition 1 emphasizes the role of demand rotation instruments. To get the intuition of how the instrument relates to the proof, consider an economy with a linear inverse demand with a demand rotation instrument and a linear marginal cost. bresnahanOligopoly1982 considers this environment. Then, assume that the true conduct is the perfect competition, that is, $\theta=0$, and the alternative conduct is the monopoly, that is, $\theta^{*}=1$. Figure 1 illustrates this environment. Note that the equilibrium quantity under perfect competition is determined by the intersection of the inverse demand function and the marginal cost function $MC$. The equilibrium quantity under the monopoly is determined by the intersection of the marginal revenue $MR^{*}$ and the marginal cost function $MC^{*}$.
[…]
(a) Step 1: Observable equivalence holds.
[…]
(b) Step 2: Demand rotation changes $D$ and $MR^{*}$
[…]
(c) Step 3.1: The intercept of $MC^{*}$ changes along with the demand rotation to keep $E=E^{*}$.
[…]
(d) Step 3.2: The slope of $MC^{*}$ changes along with the demand rotation to keep $E=E^{*}$.
Figure 1: Intuition of the role of the demand rotation instrument and identification
Note: The figures illustrate the intuition of how the demand rotation instrument works to identify the conduct parameter. $MC$ is the true marginal cost function, and $MC^{*}$ is the marginal cost function that rationalizes the monopoly conduct. Step 1 illustrates that the monopoly and the perfect competition are observationally equivalent. In step 2, the demand rotation changes the intercept and slope of the demand function without changing the equilibrium point under the perfect competition, but changes the equilibrium point under the monopoly. Step 3 illustrates that to keep the equilibrium point under the monopoly, $MC^{*}$ should change along with the demand rotation, which is impossible because the marginal cost function is independent of the demand shifter.
In Figure 1(a), the perfect competition and the monopoly lead to the same equilibrium $E$. Then, consider the change in a demand rotation instrument. In Figure 1(b), the demand rotation instrument changes the intercept and slope of the demand function without changing the equilibrium point under the true model. Under the monopoly, the new equilibrium point is $E^{*}$, which is different from $E$, and hence the observational equivalence is violated. So as to lead to the same equilibrium after the change in the demand rotation instrument ($E=E^{*})$, $MC^{*}$ should shift as in Figure 1(c) and 1(d). Note that we changed only the demand shifter, and hence the cost shifter is unchanged. Thus, for the observational equivalence to hold, the shift in $MC^{*}$ should be derived from the change in the demand rotation instrument, which implies that $MC^{*}$ is a function of the demand shifter. However, this is impossible because it violates Assumption 1, that is, the demand shifter should not affect the marginal cost function. The transformation in Lemma 2 clearly shows the possibility of the dependence of the marginal cost function on demand shifters, and hence we need to shut down the effect of demand shifters, which leads to the inverse demand function (15).
The technical problem in Lau’s proof:
To show Claim 1, Lau follows the same logic as in Lemma 1, but he obtains a slightly different equation from (18):
$\frac{\partial MC^{*}}{\partial X^{s}_{j}}(Q,\bm{X}^{s})=\mu(Q,\bm{X}^{s})\frac{\partial MC}{\partial X^{s}_{j}}(Q,\bm{X}^{s}),\quad\forall Q,\bm{X}^{s},$ (37)
where $\mu(Q,\bm{X}^{s})$ depends only on $Q$ and $\bm{X}^{s}$. This corresponds to Equation (15) in lauIdentifying1982. Then he concludes that there is a transformation $T$ such that
$MC^{*}(Q^{e},\bm{X}^{s})=T(MC(Q^{e},\bm{X}^{s}),Q^{e}).$ (38)
The existence of the transformation $T$ is shown by integrating (37) with respect to the cost shifter when it is a scalar, and by applying Lemma B.1 from goldmanNote1964 when the cost shifter is a vector. Note also that in our approach, both (17) and (18) are necessary to derive the transformation between marginal cost functions. In contrast, while Lau derives both equations (17) and (18) in his proof, he relies solely on (18) to show the non-identification.
While Lau does not explicitly define $\mu$ in his formulation, it can be verified that his $\mu$ corresponds to our $\lambda$. However, he neglects a point that the function $\mu$ implicitly depends on the demand shifter $\bm{X}^{d}$, and this dependence is not addressed in his analysis. As long as the effect of the demand shifter cannot be eliminated from $\mu$, the transformation $T$ between $MC$ and $MC^{*}$ will necessarily depend on $\bm{X}^{d}$, which violates Assumption 1. Unfortunately, Lau does not provide a justification for why the effect of the demand shifter can be removed from $\mu$, and hence it leaves a critical gap in the argument. Furthermore, Lau assumes that the inverse demand functions are twice continuously differentiable, which allows him to apply Lemma B.1 to justify the existence of the transformation $T$. To validly eliminate the effect of $\bm{X}^{d}$ from $\mu$, it is necessary to differentiate $\mu$ with respect to $\bm{X}^{d}$ and is required that the derivative is equal to zero. This step, in turn, requires that the inverse demand function be at least three-times continuously differentiable as in our proof. Therefore, the twice-continuous differentiability is not enough to characterize the non-identification.
4 Identification of All Model Components
So far, as in lauIdentifying1982, we have characterized the non-identification problem by assuming that the inverse demand function is already identified and then have not discussed the identification of the inverse demand function given the data. When the inverse demand function is also identified, all components in the conduct parameter model can be identified. Therefore, we check whether the inverse demand function, where the conduct parameter and the marginal cost function are not identified, can be identified. Unfortunately, we have the following proposition:
Proposition 3.
An inverse demand function cannot be identified when it has the following form:
$P(Q,\bm{X}^{d})=Q^{\alpha}r(\bm{X}^{d})+s(Q),$ (39)
where $\alpha$ is a constant, and $r$ and $s$ are twice continuously differentiable functions.
[…]
Another unfortunate fact is that demand rotation instruments do not help identify an inverse demand function. The following example demonstrates this point, and hence demand rotation instruments cannot be a sufficient condition for the identification of the inverse demand function:
Example 1.
Suppose the inverse demand function is given as
$P(Q,\bm{X}^{d})=Q^{\alpha}r(\bm{X}^{d})+s(Q)+r(\bm{X}^{d}).$ (40)
Here, $\bm{X}^{d}$ can work as a demand rotation instrument through $r$. However, in the same way as in Proposition 3, it is easy to see that for any $r$ and $s$, the inverse demand functions with $\tilde{r}(\bm{X}^{d})\equiv r(\bm{X}^{d})+1$ and $\tilde{s}(Q)\equiv s(Q)-Q^{\alpha}-1$ lead to the same price for any quantity and demand shifter, which implies that the inverse demand function is not identified.
Therefore, our results so far imply that the identification of the conduct parameter is guaranteed once the inverse demand function includes demand rotation instrument, but the inverse demand function with demand rotation instruments should be identified in advance.
5 Other Discussions
The results so far are true whenever we admit the conduct parameter approach. However, the conduct parameter approach has been criticized by several papers in some aspects. In this section, we discuss the micro-foundations of the conduct parameter approach and the accuracy of the conduct parameter approach.
Micro-foundations of Conduct Parameter Approach
The conduct parameter approach is based on the conjectural variation model. In the conjectural variation model, each firm has a conjecture about how the competitors will react to the firm’s action. A major problem of the conjectural variation model is that the conjecture and the reaction function do not coincide at the equilibrium. bresnahanDuopoly1981, bresnahanExistence1983 propose a Consistent Conjecture Equilibrium (CCE) which requires that the consistency between the conjecture and the reaction function is satisfied at the equilibrium. However, the existence and uniqueness of the CCE is not guaranteed in general (klempererConsistent1988, robsonExistence1983).
However, several alternative micro-foundations have been proposed. For example, escrihuela-villarNote2015 shows that the corporation coefficient approach is equivalent to the conjectural variation model in some specific setting. In this model, each firm maximizes its profit plus the weighted sum of its competitors’ profits, where the weight is called the corporation coefficient. menezesStrategic2020 investigates the supply function approach where firms choose not quantity but supply schedule based on price. menezesCompetition2023 show the relationship between the conduct parameter approach and the supply function approach in a linear demand and a linear marginal cost model case, and menezesStrategic2020 briefly discuss how to estimate the competitiveness given the information of marginal cost.
Accuracy of the Conduct Parameter Approach
While this paper only focuses on the identification problem, cortsConduct1999 criticizes that the conduct parameter approach can be inaccurate when the data generation process is out of the conduct parameter model. For example, while a repeated game model can sustain the equilibrium quantity under the joint-profit maximization, the conduct parameter estimate could be biased from $\theta=1$. At least, as cortsConduct1999 and magnolfiComparison2022 mention, the critics can be avoided by assuming the data generation process is a static competition and derived from the equilibrium condition in (1).
6 Conclusion
This paper revisited the identification of the conduct parameter in homogeneous product markets. In the literature, Lau considers the identification problem in a fairly generalized setting. While lauIdentifying1982 characterizes the non-identification condition, we point out some problems in Lau’s paper and provide a novel characterization of the non-identification condition. Based on the new characterization, we find that the non-identification is equivalent to the absence of demand shifters that work as a demand rotation instrument proposed in bresnahanOligopoly1982. We then consider the identification of all primitives in the model and show that when some demand shifters always work as demand rotation instruments, the inverse demand function cannot be identified, which immediately implies that the conduct parameter and the marginal cost function are also not identified. Therefore, while a demand rotation instrument is a key for the identification of the conduct parameter, it could prevent us from identifying it.
Acknowledgments
This paper is based on Chapter 3 of Matsumura’s dissertation submitted to Rice University. We thank Jeremy Fox for his valuable advice. We also thank the participants of Professor Fox’s group meeting for their helpful comments. We were supported by JST ERATO Grant Number JPMJER2301 and Otani was supported by JSPS 24K22604 and 25K16620.
Appendix
Appendix A Omitted Proofs
In this section, we provide the proofs of the lemmas and theorems in the main text. In what follows, we suppress the arguments of functions when no confusion arises.
A.1 Proof of Lemma 1
Assume that given $\bm{X}^{d}$ and $\bm{X}^{s}$, we can solve the equilibrium condition for the equilibrium quantity $Q^{e}$. For notational simplicity, we define
$F(Q^{e},\bm{X}^{d},\bm{X}^{s};\theta,MC)\equiv P(Q^{e},\bm{X}^{d})+\theta Q^{e}\frac{\partial P}{\partial Q}(Q^{e},\bm{X}^{d})-MC(Q^{e},\bm{X}^{s}).$ (41)
Then, by Assumption 4, we can apply the implicit function theorem, and for the reduced-form equation $Q^{e}=h_{q}(\bm{X}^{d},\bm{X}^{s})$, the gradient of the reduced-form function $h_{q}$ with respect to $\bm{X}^{d}$ and $\bm{X}^{s}$ is given by
$\nabla h_{q}(\bm{X}^{d},\bm{X}^{s})=\left(-\dfrac{\frac{\partial F}{\partial X^{m}_{i}}(h_{q}(\bm{X}^{d},\bm{X}^{s}),\bm{X}^{d},\bm{X}^{s})}{\frac{\partial F}{\partial Q}(h_{q}(\bm{X}^{d},\bm{X}^{s}),\bm{X}^{d},\bm{X}^{s})}\right)_{\begin{subarray}{c}m=d,s\\ i=1,\ldots,K_{m}\end{subarray}}.$ (42)
The derivatives of $F$ for each variable are for $i=1,\ldots,K_{d}$ and $i=1,\ldots,K_{s}$,
$\frac{\partial F}{\partial X^{d}_{i}}(Q^{e},\bm{X}^{d},\bm{X}^{s};\theta,MC)$ $=\frac{\partial P}{\partial X^{d}_{i}}(Q^{e},\bm{X}^{d})+\theta\frac{\partial^{2}P}{\partial X^{d}_{i}\partial Q}(Q^{e},\bm{X}^{d})Q^{e},$ (43)
$\frac{\partial F}{\partial X^{s}_{i}}(Q^{e},\bm{X}^{d},\bm{X}^{s};\theta,MC)$ $=-\frac{\partial MC}{\partial X^{s}_{i}}(Q^{e},\bm{X}^{s}),$ (44)
$\frac{\partial F}{\partial Q}(Q^{e},\bm{X}^{d},\bm{X}^{s};\theta,MC)$ $=(1+\theta)\frac{\partial P}{\partial Q}(Q^{e},\bm{X}^{d})+\theta Q^{e}\frac{\partial^{2}P}{\partial Q^{2}}(Q^{e},\bm{X}^{d})-\frac{\partial MC}{\partial Q}(Q^{e},\bm{X}^{s}).$ (45)
Thus, the derivative of $h$ with respect to $X^{d}_{i}$ and $X^{s}_{i}$ are given as for $i=1,\ldots,K_{d}$ and $i=1,\ldots,K_{s}$,
$\frac{\partial h_{q}}{\partial X^{d}_{i}}(\bm{X}^{d},\bm{X}^{s})=-\frac{\frac{\partial P}{\partial X^{d}_{i}}(Q^{e},\bm{X}^{d})+\theta Q^{e}\frac{\partial^{2}P}{\partial X^{d}_{i}\partial Q}(Q^{e},\bm{X}^{d})}{(1+\theta)\frac{\partial P}{\partial Q}(Q^{e},\bm{X}^{d})+\theta Q^{e}\frac{\partial^{2}P}{\partial Q^{2}}(Q^{e},\bm{X}^{d})-\frac{\partial MC}{\partial Q}(Q^{e},\bm{X}^{s})},$ (46)
and
$\frac{\partial h_{q}}{\partial X^{s}_{i}}(\bm{X}^{d},\bm{X}^{s})$ $=\frac{\frac{\partial MC}{\partial X^{s}_{i}}(Q^{e},\bm{X}^{s})}{(1+\theta)\frac{\partial P}{\partial Q}(Q^{e},\bm{X}^{d})+\theta Q^{e}\frac{\partial^{2}P}{\partial Q^{2}}(Q^{e},\bm{X}^{d})-\frac{\partial MC}{\partial Q}(Q^{e},\bm{X}^{s})}.$ (47)
Note that the same argument can be applied to the equilibrium condition (7) of the alternative model. Recall that the non-identification implies that $Q^{e}=h_{q}(\bm{X}^{d},\bm{X}^{s})=h_{q}^{*}(\bm{X}^{d},\bm{X}^{s})$ for all $\bm{X}^{d}$ and $\bm{X}^{s}$, and hence we must have
$\nabla h_{q}^{*}(\bm{X}^{d},\bm{X}^{s})=\nabla h_{q}(\bm{X}^{d},\bm{X}^{s})\quad\forall\bm{X}^{d},\bm{X}^{s}.$ (48)
From (42), this implies that for all $Q^{e}$, $\bm{X}^{d}$, and $\bm{X}^{s}$,
$\begin{pmatrix}\frac{\partial F}{\partial\bm{X}^{d}_{1}}(Q^{e},\bm{X}^{d},\bm{X}^{s};\theta^{*},MC^{*})\\ \vdots\\ \frac{\partial F}{\partial\bm{X}^{d}_{K_{d}}}(Q^{e},\bm{X}^{d},\bm{X}^{s};\theta^{*},MC^{*})\\ \frac{\partial F}{\partial X^{s}_{1}}(Q^{e},\bm{X}^{d},\bm{X}^{s};\theta^{*},MC^{*})\\ \vdots\\ \frac{\partial F}{\partial X^{s}_{K_{s}}}(Q^{e},\bm{X}^{d},\bm{X}^{s};\theta^{*},MC^{*})\end{pmatrix}=\lambda(Q^{e},\bm{X}^{d},\bm{X}^{s})\begin{pmatrix}\frac{\partial F}{\partial\bm{X}^{d}_{1}}(Q^{e},\bm{X}^{d},\bm{X}^{s};\theta,MC)\\ \vdots\\ \frac{\partial F}{\partial\bm{X}^{d}_{K_{d}}}(Q^{e},\bm{X}^{d},\bm{X}^{s};\theta,MC)\\ \frac{\partial F}{\partial X^{s}_{1}}(Q^{e},\bm{X}^{d},\bm{X}^{s};\theta,MC)\\ \vdots\\ \frac{\partial F}{\partial X^{s}_{K_{s}}}(Q^{e},\bm{X}^{d},\bm{X}^{s};\theta,MC)\end{pmatrix}$, (49)
where $\lambda(Q^{e},\bm{X}^{d},\bm{X}^{s})$ is defined as
$\lambda(Q^{e},\bm{X}^{d},\bm{X}^{s})\equiv\frac{(1+\theta^{*})\frac{\partial P}{\partial Q}(Q^{e},\bm{X}^{d})+\theta^{*}Q^{e}\frac{\partial^{2}P}{\partial Q^{2}}(Q^{e},\bm{X}^{d})-\frac{\partial MC^{*}}{\partial Q}(Q^{e},\bm{X}^{s})}{(1+\theta)\frac{\partial P}{\partial Q}(Q^{e},\bm{X}^{d})+\theta Q^{e}\frac{\partial^{2}P}{\partial Q^{2}}(Q^{e},\bm{X}^{d})-\frac{\partial MC}{\partial Q}(Q^{e},\bm{X}^{s})}.$ (50)
Assumption 4 guarantees that $\lambda(Q^{e},\bm{X}^{d},\bm{X}^{s})$ is nonzero and finite. From the first $K_{d}$ rows in (49), (17) and for the later $K_{s}$ rows, (18) hold.
A.2 Proof of Lemma 2
From (17), we have (we suppress the arguments of the functions for brevity)
$\left[\frac{\partial P}{\partial X^{d}_{i}}+\theta^{*}Q^{e}\frac{\partial^{2}P}{\partial X^{d}_{i}\partial Q}\right]\left[(1+\theta)\frac{\partial P}{\partial Q}+\theta Q^{e}\frac{\partial^{2}P}{\partial Q^{2}}-\frac{\partial MC}{\partial Q}\right]$ (51)
$\hskip 113.81102pt=\left[\frac{\partial P}{\partial X^{d}_{i}}+\theta Q^{e}\frac{\partial^{2}P}{\partial X^{d}_{i}\partial Q}\right]\left[(1+\theta^{*})\frac{\partial P}{\partial Q}+\theta^{*}Q^{e}\frac{\partial^{2}P}{\partial Q^{2}}-\frac{\partial MC^{*}}{\partial Q}\right]$ (52)
$\frac{\frac{\partial P}{\partial X^{d}_{i}}+\theta^{*}Q\frac{\partial^{2}P}{\partial X^{d}_{i}\partial Q}}{\frac{\partial P}{\partial X^{d}_{i}}+\theta Q\frac{\partial^{2}P}{\partial X^{d}_{i}\partial Q}}\left[(1+\theta)\frac{\partial P}{\partial Q}+\theta Q^{e}\frac{\partial^{2}P}{\partial Q^{2}}-\frac{\partial MC}{\partial Q}\right]$ (53)
$\hskip 113.81102pt=\left[(1+\theta^{*})\frac{\partial P}{\partial Q}+\theta^{*}Q^{e}\frac{\partial^{2}P}{\partial Q^{2}}-\frac{\partial MC^{*}}{\partial Q}\right]$ (54)
$\frac{\partial MC^{*}}{\partial Q}=\underbrace{(1+\theta^{*})\frac{\partial P}{\partial Q}+\theta^{*}Q^{e}\frac{\partial^{2}P}{\partial Q^{2}}-C_{i}(Q,\bm{X}^{d})\left[(1+\theta)\frac{\partial P}{\partial Q}+\theta Q^{e}\frac{\partial^{2}P}{\partial Q^{2}}\right]}_{(\ast)}+C_{i}(Q,\bm{X}^{d})\frac{\partial MC}{\partial Q}.$ (55)
The terms in $(\ast)$ can be more simplified as
$(1+\theta^{*})\frac{\partial P}{\partial Q}+\theta^{*}Q\frac{\partial^{2}P}{\partial Q^{2}}-\frac{\frac{\partial P}{\partial X^{d}_{i}}+\theta^{*}Q\frac{\partial^{2}P}{\partial X^{d}_{i}\partial Q}}{\frac{\partial P}{\partial X^{d}_{i}}+\theta Q\frac{\partial^{2}P}{\partial X^{d}_{i}\partial Q}}\left[(1+\theta)\frac{\partial P}{\partial Q}+\theta Q\frac{\partial^{2}P}{\partial Q^{2}}\right]$ (56)
$=\frac{1}{\frac{\partial P}{\partial X^{d}_{i}}+\theta\frac{\partial^{2}P}{\partial X^{d}_{i}\partial Q}Q}\Bigg[\left((1+\theta^{*})\frac{\partial P}{\partial Q}+\theta^{*}Q\frac{\partial^{2}P}{\partial Q^{2}}\right)\left(\frac{\partial P}{\partial X^{d}_{i}}+\theta Q\frac{\partial^{2}P}{\partial X^{d}_{i}\partial Q}\right)$ (57)
$\hskip 113.81102pt-\left((1+\theta)\frac{\partial P}{\partial Q}+\theta Q\frac{\partial^{2}P}{\partial Q^{2}}\right)\left(\frac{\partial P}{\partial X^{d}_{i}}+\theta^{*}Q\frac{\partial^{2}P}{\partial X^{d}_{i}\partial Q}\right)\Bigg]$ (58)
$=\frac{\theta^{*}-\theta}{\frac{\partial P}{\partial X^{d}_{i}}+\theta\frac{\partial^{2}P}{\partial X^{d}_{i}\partial Q}Q}\left[\frac{\partial P}{\partial Q}\frac{\partial P}{\partial X^{d}_{i}}+Q\frac{\partial P}{\partial X^{d}_{i}}\frac{\partial^{2}P}{\partial^{2}Q}-Q\frac{\partial P}{\partial Q}\frac{\partial^{2}P}{\partial X^{d}_{i}\partial Q}\right],$ (59)
which corresponds to $D_{i}(Q,\bm{X}^{d})$ in (24), and hence we have
$\frac{\partial MC^{*}}{\partial Q}(Q,\bm{X}^{s})=D_{i}(Q,\bm{X}^{d})+C_{i}(Q,\bm{X}^{d})\frac{\partial MC}{\partial Q}(Q,\bm{X}^{s}).$ (60)
Next, (17) and (18) imply that for $i=1,\ldots,K_{d}$ and $j=1,\ldots,K_{s}$,
$\lambda(Q^{e},\bm{X}^{d},\bm{X}^{s})=\frac{\frac{\partial MC^{*}}{\partial X^{s}_{j}}(Q^{e},\bm{X}^{s})}{\frac{\partial MC}{\partial X^{s}_{j}}(Q^{e},\bm{X}^{s})}=\frac{\frac{\partial P}{\partial X^{d}_{i}}(Q^{e},\bm{X}^{d})+\theta^{*}Q^{e}\frac{\partial^{2}P}{\partial X^{d}_{i}\partial Q}(Q^{e},\bm{X}^{d})}{\frac{\partial P}{\partial X^{d}_{i}}(Q^{e},\bm{X}^{d})+\theta Q^{e}\frac{\partial^{2}P}{\partial X^{d}_{i}\partial Q}(Q^{e},\bm{X}^{d})}\equiv C_{i}(Q,\bm{X}^{d}),$ (61)
Then, we have
$\frac{\partial MC^{*}}{\partial X^{s}_{j}}(Q,\bm{X}^{s})=C_{i}(Q,\bm{X}^{d})\frac{\partial MC}{\partial X^{s}_{j}}(Q,\bm{X}^{s}).$ (62)
A.3 Proof of Lemma 3
Note that when $\theta=0$, (25) implies that $\frac{\partial P}{\partial X^{d}_{i}}(Q,\bm{X}^{d})=0$ for all $i\in\mathcal{I}$. This cannot be held because it violates Assumption 2. Therefore, we must assume that $\theta\neq 0$ for the identification.
Let $u_{i}(Q,\bm{X}^{d})\equiv\frac{\partial P}{\partial X^{d}_{i}}(Q,\bm{X}^{d})$. Again, due to Assumption 2, we have that $u_{i}(Q,\bm{X}^{d})\neq 0$ for all $Q$ and $\bm{X}^{d}$. Rewrite (25) for $i\in\mathcal{I}$ as
$\frac{\frac{\partial u}{\partial Q}(Q,\bm{X}^{d})}{u_{i}(Q,\bm{X}^{d})}=-\frac{1}{\theta Q}\Longrightarrow\frac{\partial}{\partial Q}\log|u_{i}(Q,\bm{X}^{d})|=-\frac{1}{\theta Q}.$ (63)
By integrating both sides with respect to $Q$, we have
$\log|u_{i}(Q,\bm{X}^{d})|=-\frac{1}{\theta}\log|Q|+R_{i}(\bm{X}^{d}),$ (64)
where $R_{i}(\bm{X}^{d})$ is a function of $\bm{X}^{d}$. By taking the exponential of both sides, we have
$|u_{i}(Q,\bm{X}^{d})|=|Q|^{-\frac{1}{\theta}}r_{i}(\bm{X}^{d}),$ (65)
where $r_{i}(\bm{X}^{d})=\exp(R_{i}(\bm{X}^{d}))$. Note that we can remove the absolute value for $Q$ because it can be assumed that $Q\geq 0$. By removing the absolute value operator for $u_{i}$, we have two solutions
$u_{i}(Q,\bm{X}^{d})=\pm Q^{-\frac{1}{\theta}}r_{i}(\bm{X}^{d}).$ (66)
Because $r_{i}$ is an arbitrary function, we can unify these solutions and can simply put the solution as
$u_{i}(Q,\bm{X}^{d})=Q^{-\frac{1}{\theta}}r_{i}(\bm{X}^{d}).$ (67)
This implies that the derivative of $P$ with respect to $X^{d}_{i}$ is a separable function of $Q$ and $\bm{X}^{d}$. Hence, it is natural to think that $r_{i}(\bm{X}^{d})$ is a derivative of a function of $\bm{X}^{d}$ with respect to $X^{d}_{i}$. In other words, there is a function $r(\bm{X}^{d})$ of $\bm{X}^{d}$ such that $r_{i}(\bm{X}^{d})=\frac{\partial r(\bm{X}^{d})}{\partial X^{d}_{i}}$ holds. Therefore, we have for $i\in\mathcal{I}$,
$u_{i}(Q,\bm{X}^{d})=\frac{\partial r(\bm{X}^{d})}{\partial X^{d}_{i}}Q^{-\frac{1}{\theta}}.$ (68)
By integrating both sides with respect to $X^{d}_{i}$, we have
$P(Q,\bm{X}^{d})=Q^{-\frac{1}{\theta}}r(\bm{X}^{d})+s_{i}(Q,\bm{X}^{d}_{-i}),\quad i\in\mathcal{I},$ (69)
where $s_{i}(Q,\bm{X}^{d}_{-i})$ is a function of $Q$ and $\bm{X}^{d}_{-i}$ where $\bm{X}^{d}_{-i}$ is the vector of $\bm{X}^{d}$ excluding $X^{d}_{i}$. To meet Assumption 3, we must have that $R(\bm{X}^{d})$ and $s_{i}(Q,\bm{X}^{d}_{-i})$ are at least twice-continuously differentiable.
When $\bm{X}^{d}$ is a scalar or when $\bm{X}^{d}$ is a vector but $|\mathcal{I}|\geq K_{d}-1$, there is no argument in $\bm{X}^{d}_{-i}$, and hence we can remove the index of $i$ from the function $s_{i}$. Therefore, $s(Q,\bm{X}^{d}_{-i})=s(Q)$ holds, and hence we have (27). When $\bm{X}^{d}$ is a vector and $K_{d}-2\geq|\mathcal{I}|\geq 2$, we can remove the index $i$ from $s_{i}$ in the following way. Pick up any $i$ and $j$ in $\mathcal{I}$ such that $i\neq j$, and then the derivative of (69) with respect to $X^{d}_{j}$ is given by
$\frac{\partial P}{\partial X^{d}_{j}}(Q,\bm{X}^{d})=\frac{\partial r(\bm{X}^{d})}{\partial X^{d}_{j}}Q^{-\frac{1}{\theta}}+\frac{\partial s_{i}(Q,\bm{X}^{d}_{-i})}{\partial X^{d}_{j}},$ (70)
and
$\frac{\partial P}{\partial X^{d}_{j}}(Q,\bm{X}^{d})=\frac{\partial r(\bm{X}^{d})}{\partial X^{d}_{j}}Q^{-\frac{1}{\theta}}.$ (71)
Note that $s_{j}$ is a function of the demand shifters without $X_{j}$, we don’t have the term relating to $s_{j}$ in the second equation. Therefore, by comparing the above two equations, we can conclude that the second term in the first equation vanishes:
$\frac{\partial s_{i}(Q,\bm{X}^{d}_{-i})}{\partial X^{d}_{j}}=0,$ (72)
which implies that $s_{i}(Q,\bm{X}^{d}_{-i})$ is independent of $X^{d}_{j}$. By applying the same argument to all $i\in\mathcal{I}$, we can show that $s_{i}(Q,\bm{X}^{d}_{-i})$ is independent of the demand shifters whose indices are in $\mathcal{I}$. That is, we can reduce $s_{i}(Q,\bm{X}^{d}_{-i})$ to $s_{i}(Q,\bm{X}^{d}_{-\mathcal{I}})$ where $\bm{X}^{d}_{-\mathcal{I}}$ is the vector of $\bm{X}^{d}$ excluding the demand shifters whose indices are in $\mathcal{I}$. Then, by comparing (69) for all $i\in\mathcal{I}$, we have
$s_{i}(Q,\bm{X}^{d}_{-\mathcal{I}})=s_{j}(Q,\bm{X}^{d}_{-\mathcal{I}}),$ (73)
for all $i,j\in\mathcal{I}$. This can leads to the symmetry of $s_{i}$ for all $i\in\mathcal{I}$, which implies that $s(Q,\bm{X}^{d}_{-\mathcal{I}})=s_{i}(Q,\bm{X}^{d}_{-\mathcal{I}})$ for all $i\in\mathcal{I}$. Therefore, we can write the inverse demand function as
$P(Q,\bm{X}^{d})=Q^{-\frac{1}{\theta}}r(\bm{X}^{d})+s(Q,\bm{X}^{d}_{-\mathcal{I}}).$ (74)
When $|\mathcal{I}|=1$, we have $\bm{X}^{d}_{-\mathcal{I}}=\bm{X}^{d}_{-i}$, and hence we can keep the index for $s_{i}$, but it is also fine to write it as $s(Q,\bm{X}^{d}_{-i})=s_{i}(Q,\bm{X}^{d}_{-i})$.
A.4 Proof of Proposition 1
Because $C_{i}(Q,\bm{X}^{d})$ and $D_{i}(Q,\bm{X}^{d})$ should be independent of $\bm{X}^{d}$ simultaneously, we can first characterize the class of inverse demand functions that make $C_{i}(Q,\bm{X}^{d})$ independent of $\bm{X}^{d}$. Then, we characterize the class of inverse demand functions that make $D_{i}(Q,\bm{X}^{d})$ independent of $\bm{X}^{d}$ by substituting the derived inverse demand function into $D_{i}(Q,\bm{X}^{d})$.
Step 1: Necessity for $C_{i}$ is independent of $\bm{X}^{d}$
When $C_{i}(Q,\bm{X}^{d})$ is independent of $\bm{X}^{d}$, the derivative of $C_{i}(Q,\bm{X}^{d})$ with respect to $\bm{X}^{d}$ is zero. Let $u_{i}(Q,\bm{X}^{d})\equiv\frac{\partial P}{\partial X^{d}_{i}}(Q,\bm{X}^{d})$. Due to Assumption 2, we have that $u_{i}(Q,\bm{X}^{d})\neq 0$ for all $Q$ and $\bm{X}^{d}$. Then, the derivative of $C_{i}$ with respect to $X^{d}_{j}$ for $j=1,\ldots,K_{d}$ is given by
$\frac{\partial C_{i}}{\partial X^{d}_{j}}(Q,\bm{X}^{d})$ $=\frac{(\theta^{*}-\theta)Q}{\left(u_{i}+\theta Q\frac{\partial u_{i}}{\partial Q}\right)^{2}}\left[\frac{\partial u_{i}}{\partial X^{d}_{j}\partial Q}u_{i}-\frac{\partial u_{i}}{\partial Q}\frac{\partial u_{i}}{\partial X^{d}_{j}}\right].$ (75)
Note that because (25) does not hold, the denominator is not zero. As $\theta\neq\theta^{*}$, the derivative becomes zero when (1) $Q=0$ or (2) the term in the bracket is zero. Because the former case is trivial, we focus on the latter case. In this case, we have
$\frac{\partial u_{i}}{\partial X^{d}_{j}\partial Q}u_{i}-\frac{\partial u_{i}}{\partial Q}\frac{\partial u_{i}}{\partial X^{d}_{j}}=0.$ (76)
Now, instead of using (76) directly, we use the following relationship:
$\frac{\partial^{2}}{\partial X^{d}_{j}\partial Q}\log|u_{i}(Q,\bm{X}^{d})|$ $=\frac{\partial}{\partial Q}\left(\frac{1}{u_{i}}\frac{\partial u_{i}}{\partial X^{d}_{j}}\right)=\frac{1}{u_{i}^{2}}\left(u_{i}\frac{\partial^{2}u_{i}}{\partial X^{d}_{j}\partial Q}-\frac{\partial u_{i}}{\partial X^{d}_{j}}\frac{\partial u_{i}}{\partial Q}\right).$ (77)
Note that the assumption $u_{i}\neq 0$ implies that the inside of the log function is not zero and the denominator in the last equality is not zero. Therefore, to check the independence of $C_{i}$ with respect to $\bm{X}^{d}$, it is sufficient to solve the following partial differential equation such that
$\frac{\partial^{2}}{\partial X^{d}_{j}\partial Q}\log|u_{i}(Q,\bm{X}^{d})|=0,\quad i,j=1,\ldots,K_{d}.$ (78)
This implies that $\frac{\partial}{\partial Q}\log|u_{i}(Q,\bm{X}^{d})|$ is independent of any element in $\bm{X}^{d}$, and hence we have a function $G(Q)$ such that
$\frac{\partial}{\partial Q}\log|u_{i}(Q,\bm{X}^{d})|=G(Q).$ (79)
By integrating both sides with respect to $Q$, we have for $i=1,\ldots,K_{d}$,
$\log|u_{i}(Q,\bm{X}^{d})|=\int|G(Q)|dQ+R_{i}(\bm{X}^{d}),$ (80)
where the last term is a function that is an analogue of the constant of integration. By taking exponential both side, we have for $i=1,\ldots,K_{d}$,
$|u_{i}(Q,\bm{X}^{d})|=g(Q)r_{i}(\bm{X}^{d})$ (81)
where $r_{i}(\bm{X}^{d})\equiv\exp(R_{i}(\bm{X}^{d}))$ and $g(Q)=\exp\left(\int|G(Q)|dQ\right)$. This has two solutions, but again, because $r_{i}$ is an arbitrary function, we can unify these solutions and put the solution as
$u_{i}(Q,\bm{X}^{d})=g(Q)r_{i}(\bm{X}^{d}).$ (82)
Because $u_{i}$ is a separable function of $Q$ and $\bm{X}^{d}$, it is natural to think that $r_{i}(\bm{X}^{d})$ is a derivative of a function of $\bm{X}^{d}$ with respect to $X^{d}_{i}$. In other words, there is a function $r(\bm{X}^{d})$ of $\bm{X}^{d}$ such that $r_{i}(\bm{X}^{d})=\frac{\partial r(\bm{X}^{d})}{\partial X^{d}_{i}}$ holds. Therefore, we have
$u_{i}(Q,\bm{X}^{d})=\frac{\partial r(\bm{X}^{d})}{\partial X^{d}_{i}}g(Q).$ (83)
By taking the integration with respect to $X^{d}_{i}$ on both sides, we have
$P(Q,\bm{X}^{d})=g(Q)r(\bm{X}^{d})+s_{i}(Q,\bm{X}^{d}_{-i}),\quad i=1,\ldots,K_{d},$ (84)
where $s_{i}(Q,\bm{X}^{d}_{-i})$ is an arbitrary function of $Q$ and $\bm{X}^{d}_{-i}$ where $\bm{X}^{d}_{-i}$ is the vector of $\bm{X}^{d}$ excluding $X^{d}_{i}$. To meet Assumption 3, we must have that $R(\bm{X}^{d})$ and $s_{i}(Q,\bm{X}^{d}_{-i})$ are at least twice-continuously differentiable and $g(Q)$ is at least continuously differentiable.
We can remove the index of $s_{i}$ in the following way. Pick up any $i$, and then the derivative of (84) with respect to $X^{d}_{i}$ is given by
$\frac{\partial P}{\partial X^{d}_{i}}(Q,\bm{X}^{d})=\frac{\partial r(\bm{X}^{d})}{\partial X^{d}_{i}}g(Q)+\frac{\partial s_{i}(Q,\bm{X}^{d}_{-i})}{\partial X^{d}_{i}}=\frac{\partial r(\bm{X}^{d})}{\partial X^{d}_{i}}g(Q),$ (85)
and for any other $j\neq i$, the same derivative is given by
$\frac{\partial P}{\partial X^{d}_{i}}(Q,\bm{X}^{d})=\frac{\partial r(\bm{X}^{d})}{\partial X^{d}_{i}}g(Q)+\frac{\partial s_{j}(Q,\bm{X}^{d}_{-j})}{\partial X^{d}_{i}}.$ (86)
These imply that
$\frac{\partial s_{j}(Q,\bm{X}^{d}_{-j})}{\partial X^{d}_{i}}=0\quad\text{for all }j\neq i.$ (87)
Therefore, $s_{j}(Q,\bm{X}^{d}_{-j})$ is independent of $X^{d}_{i}$, and hence $s_{j}$ is independent of $X^{d}_{j}$ and $X^{d}_{i}$. By applying the same argument to all $i=1,\ldots,K_{d}$, we can show that $s_{i}(Q,\bm{X}^{d}_{-i})$ is independent of all elements in $\bm{X}^{d}_{-i}$, that is, $s_{i}(Q,\bm{X}^{d}_{-i})=s_{i}(Q)$ for all $i=1,\ldots,K_{d}$.
By comparing (84) for all $i=1,\ldots,K_{d}$, it is easy to see that $s_{i}(Q)=s_{j}(Q)$ for all $i,j=1,\ldots,K_{d}$. Therefore, we have a symmetry of $s_{i}$ for all $i=1,\ldots,K_{d}$, that is, $s_{i}(Q)=s(Q)$ for all $i=1,\ldots,K_{d}$. Thus, $P(Q,\bm{X}^{d})$ must be of the form
$P(Q,\bm{X}^{d})=g(Q)r(\bm{X}^{d})+s(Q).$ (88)
Step 2: Necessity for $D_{i}$ is independent of $\bm{X}^{d}$
Next, given the inverse demand function (88), we further specify the form of the inverse demand function based on $D_{i}(Q,\bm{X}^{d})$. By substituting (88) into (24), we have
$D_{i}(Q,\bm{X}^{d})=$ $\frac{\theta^{*}-\theta}{r_{i}(\bm{X}^{d})(g(Q)+\theta Qg^{\prime}(Q))}\Big[(g^{\prime}(Q)r(\bm{X}^{d})+s^{\prime}(Q))g(Q)r_{i}(\bm{X}^{d})$ (89)
$\hskip 128.0374pt+Q(g^{\prime\prime}(Q)r(\bm{X}^{d})+s^{\prime\prime}(Q))g(Q)r_{i}(\bm{X}^{d})$ (90)
$\hskip 136.5733pt-Q(g^{\prime}(Q)r(\bm{X}^{d})+s^{\prime}(Q))g^{\prime}(Q)r_{i}(\bm{X}^{d})\Big]$ (91)
$=$ $\frac{\theta^{*}-\theta}{g(Q)+\theta Qg^{\prime}(Q)}\Big[r(\bm{X}^{d})[g^{\prime}(Q)g(Q)+Qg{{}^{\prime\prime}}(Q)g(Q)-Q(g^{\prime}(Q))^{2}]$ (92)
$\hskip 99.58464pt+s^{\prime}(Q)g(Q)+Qs^{\prime\prime}(Q)g(Q)g^{\prime\prime}(Q)-Qs^{\prime}(Q)g^{\prime}(Q)\Big].$ (93)
The dependence of $D_{i}(Q,\bm{X}^{d})$ on $\bm{X}^{d}$ comes only from the first term in the bracket. Therefore, $D_{i}(Q,\bm{X}^{d})$ is independent of $\bm{X}^{d}$ if and only if
$\frac{\partial D_{i}}{\partial X^{d}_{i}}(Q,\bm{X}^{d})=\frac{(\theta^{*}-\theta)r_{i}(\bm{X}^{d})}{g(Q)+\theta Qg^{\prime}(Q)}[g^{\prime}(Q)g(Q)+Qg{{}^{\prime\prime}}(Q)g(Q)-Q(g^{\prime}(Q))^{2}]=0.$ (94)
As Assumption 2 implies that $r_{i}(\bm{X}^{d})$ is nonzero, we must have the terms in the bracket to be zero:
$g^{\prime}(Q)g(Q)+Qg{{}^{\prime\prime}}(Q)g(Q)-Q(g^{\prime}(Q))^{2}=0.$ (95)
Let $v(Q)=\frac{g^{\prime}(Q)}{g(Q)}$. Because $v^{\prime}(Q)=\frac{g^{\prime\prime}(Q)g(Q)-(g^{\prime}(Q))^{2}}{g(Q)^{2}}$, dividing (95) by $g(Q)^{2}$ gives
$v(Q)+Qv^{\prime}(Q)=0.$ (96)
This is a first-order linear differential equation. To solve this differential equation, we consider two cases.
Case 1: $v(Q)=0$ for all $Q$.
This happens when $g(Q)$ is a constant. In this case, (95) holds immediately, and hence (96) also holds. Let $g(Q)=C$ for some nonzero constant $C$. Then, we have
$P(Q,\bm{X}^{d})=Cr(\bm{X}^{d})+s(Q).$ (97)
For simplicity, we can absorb the constant $C$ into $r(\bm{X}^{d})$ and $s(Q)$, and hence we have
$P(Q,\bm{X}^{d})=r(\bm{X}^{d})+s(Q).$ (98)
Case 2: $v(Q)\neq 0$ for all $Q$.
In this case, (96) implies that
$\frac{v^{\prime}(Q)}{v(Q)}=-\frac{1}{Q}.$ (99)
Then, it can be written as
$\frac{d}{dQ}\log|v(Q)|=-\frac{d}{dQ}\log|Q|.$ (100)
Integrating both sides with respect to $Q$, we have
$\log|v(Q)|=-\log|Q|+a_{1},$ (101)
where $a_{1}\in\mathbb{R}$ is a constant. By taking the exponential of both sides, we have
$|v(Q)|=\frac{\alpha}{|Q|},$ (102)
where $\alpha=\exp(a_{1})$ is a positive constant. This has two solutions
$v(Q)=\frac{g^{\prime}(Q)}{g(Q)}=\pm\frac{\alpha}{|Q|}.$ (103)
[…]
$\frac{d}{dQ}\log|g(Q)|=\alpha\frac{d}{dQ}\log|Q|.$ (104)
Again, by integrating both sides with respect to $Q$, we have
$\log|g(Q)|=\alpha\log|Q|+a_{2},$ (105)
where $a_{2}\in\mathbb{R}$ is a constant. By taking the exponential of both sides, we have
$|g(Q)|=\alpha_{2}|Q|^{\alpha},$ (106)
where $\alpha_{2}=\exp(a_{2})$ is a positive constant. Again, this has two solutions
$g(Q)=\pm\alpha_{2}|Q|^{\alpha}.$ (107)
[…]
$P(Q,\bm{X}^{d})=Q^{\alpha}r(\bm{X}^{d})+s(Q).$ (108)
Here, $\alpha_{2}$ is absorbed into $r$.
While $\alpha$ is nonzero in (108), by allowing that $\alpha=0$, it can include (98) as a special case. At the same time, recall that we require that (25) does not hold for any $i$. However, when $\alpha=-\frac{1}{\theta}$, we have the inverse demand function is equal to (27) where $\mathcal{I}=\emptyset$, which implies that the conduct parameter and the marginal cost function are identified. Therefore, we should have $\alpha\neq-\frac{1}{\theta}$ to ensure the non-identification of the conduct parameter. This completes the proof.
A.5 Proof of Proposition 2
From the proof of Proposition 1, we know that $C_{i}$ and $D_{i}$ are independent of $\bm{X}^{d}$ under (15). Under the inverse demand function (88), we have
$\hat{C}_{i}(Q,\bm{X}^{d})=\frac{Q^{\alpha}h^{\prime}(\bm{X}^{d})+\theta^{*}\alpha Q^{\alpha}h^{\prime}(\bm{X}^{d})}{Q^{\alpha}h^{\prime}(\bm{X}^{d})+\theta\alpha Q^{\alpha}h^{\prime}(\bm{X}^{d})}=\frac{1+\theta^{*}\alpha}{1+\theta\alpha}.$ (109)
and
$\hat{D}_{i}(Q,\bm{X}^{d})$ $=\frac{\theta^{*}-\theta}{Q^{\alpha}+\theta\alpha Q^{\alpha}}\left[s^{\prime}(Q)Q^{\alpha}+Qk^{\prime\prime}(Q)Q^{\alpha}-s^{\prime}(Q)\alpha Q^{\alpha}\right]$ (110)
$=\frac{\theta^{*}-\theta}{1+\theta\alpha}\left[\frac{d}{dQ}Qs^{\prime}(Q)-\alpha s^{\prime}(Q)\right].$ (111)
Then, consider a transformation of the derivative of a function $f(Q,X)$ with respect to $Q$ based on (20) such that
$T_{Q}\left(f(Q,X),Q\right)$ $\equiv\hat{D}_{i}(Q,\bm{X}^{d})+\hat{C}_{i}(Q,\bm{X}^{d})\frac{\partial f}{\partial Q}(Q,X)$ (112)
$=\frac{\theta^{*}-\theta}{1+\theta\alpha}\left[\frac{d}{dQ}Qs^{\prime}(Q)-\alpha s^{\prime}(Q)\right]+\frac{1+\theta^{*}\alpha}{1+\theta\alpha}\frac{\partial f}{\partial Q}(Q,X).$ (113)
This can be integrated with respect to $Q$, and hence we obtain a transformation of $f(Q,X)$ as
$T\left(f(Q,X),Q\right)\equiv\frac{\theta^{*}-\theta}{1+\theta\alpha}\left[Qs^{\prime}(Q)-\alpha s(Q)\right]+\frac{1+\theta^{*}\alpha}{1+\theta\alpha}f(Q,X).$ (114)
Note that we assume the integral constant is zero for simplicity.
Define a marginal cost function $MC^{*}$ as
$MC^{*}(Q,\bm{X}^{s})\equiv\frac{\theta^{*}-\theta}{1+\theta\alpha}\left[Qs^{\prime}(Q)-\alpha s(Q)\right]+\frac{1+\theta^{*}\alpha}{1+\theta\alpha}MC(Q,\bm{X}^{s}).$ (115)
Then, by substituting the marginal revenue under $\theta$ into $T$, we can obtain the marginal revenue under $\theta^{*}$:
$\frac{\theta^{*}-\theta}{1+\theta\alpha}\left[Qs^{\prime}(Q)-\alpha s(Q)\right]+\frac{1+\theta^{*}\alpha}{1+\theta\alpha}\left[(1+\theta\alpha)Q^{\alpha}r(\bm{X}^{d})+s(Q)+\theta Qs^{\prime}(Q)\right]$ (116)
$=$ $(1+\theta^{*}\alpha)Q^{\alpha}r(\bm{X}^{d})+\frac{(\theta^{*}-\theta+(1+\theta^{*}\alpha)\theta)Qs^{\prime}(Q)+(1+\theta^{*}\alpha-(\theta^{*}-\theta)\alpha)s(Q)}{1+\theta\alpha}$ (117)
$=$ $(1+\theta^{*}\alpha)Q^{\alpha}r(\bm{X}^{d})+\frac{(1+\theta\alpha)s(Q)+\theta^{*}(1+\theta\alpha)Qs^{\prime}(Q)}{1+\theta\alpha}$ (118)
$=$ $(1+\theta^{*}\alpha)Q^{\alpha}r(\bm{X}^{d})+s(Q)+\theta^{*}Qs^{\prime}(Q).$ (119)
Then, suppose that $Q^{e}$ satisfies the equilibrium condition under $(\theta,MC)$. Then, by using the definition of $MC^{*}$ and the above observation on the marginal revenue, we can confirm that $Q^{e}$ also satisfies the equilibrium condition under $(\theta^{*},MC^{*})$:
$P(Q^{e},\bm{X}^{d})+\theta^{*}Q^{e}\frac{\partial P}{\partial Q}(Q^{e},\bm{X}^{d})-MC^{*}(Q^{e},\bm{X}^{s})$ (120)
$=$ $\frac{\theta^{*}-\theta}{1+\theta\alpha}\left[Qs^{\prime}(Q)-\alpha s(Q)\right]+\frac{1+\theta^{*}\alpha}{1+\theta\alpha}\left[(1+\theta\alpha)Q^{\alpha}r(\bm{X}^{d})+s(Q)+\theta Qs^{\prime}(Q)\right]$ (121)
$\hskip 28.45274pt-\frac{\theta^{*}-\theta}{1+\theta\alpha}\left[Qs^{\prime}(Q)-\alpha s(Q)\right]-\frac{1+\theta^{*}\alpha}{1+\theta\alpha}MC(Q^{e},\bm{X}^{s})$ (122)
$=$ $\frac{1+\theta^{*}\alpha}{1+\theta\alpha}\left[(1+\theta\alpha)Q^{\alpha}r(\bm{X}^{d})+s(Q)+\theta^{*}Qs^{\prime}(Q)-MC(Q^{e},\bm{X}^{s})\right]=0.$ (123)
The last inequality holds because $Q^{e}$ satisfies the equilibrium condition under $(\theta,MC)$. Therefore, the two models are observationally equivalent, which implies that the non-identification holds.
Appendix B Summary of Goldman and Uzawa (1964)
Let $n$ be the number of variables and $x=(x_{1},\ldots,x_{n})$ be a vector of $n$ variables. Consider a partition of $X$ into $K$ parts, $\{x^{1},\ldots,x^{K}\}$ such that $X=\bigcup_{k=1}^{K}x^{k}$ and $x^{k}\cap x^{l}=\emptyset$ for $k\neq l$.
Definition B.1.
A function is weakly separable with respect to the partition if
$\frac{\partial}{\partial x_{l}}\left(\frac{\frac{\partial f}{\partial x_{i}}(x^{1},\ldots,x^{K})}{\frac{\partial f}{\partial x_{j}}(x^{1},\ldots,x^{K})}\right)=0,\quad i,j\in x^{k},l\notin x^{k}.$ (124)
This implies that the ratio of the derivative with respect to $x_{i}$ and $x_{j}$, which are in the same category, is not affected by the change in the variables in other partitions. Intuitively, by taking the ratio, the component of $f$ relating to $x_{l}$ is canceled out, and hence the derivative of the ratio with respect to $x_{l}$ becomes zero. When $f$ is a utility function, this implies that the marginal rate of substitution between commodity $i$ and $j$ in the same partition is independent of the quantities of commodities outside $x^{k}$.
Then, goldmanNote1964 specifies the functional form that a weak separable function should satisfy.
Theorem B.1.
A function $f(x)$ is weakly separable with respect to a partition $\{x^{1},...,x^{K}\}$ if and only if $f(x)$ is of the form:
$f(X)=\Phi(r^{1}(x^{1}),\ldots,r^{K}(x^{K}))$ (125)
where $\Phi(r^{1},\ldots,r^{K})$ is a function of $K$ variables and, for each $k$, $r^{k}(x^{k})$ is a function subvector $x^{k}$ alone.
The next lemma is a key lemma in Lau’s proof.
Lemma B.1.
Let $f(x)$ and $g(x)$ be two continuously twice-differentiable functions of $n$ variables $x=(x_{1},\dots,x_{n})$. If each indifference surface is connected, and if there exists a function $\lambda(x)$ such that
$\frac{\partial f}{\partial x_{i}}(x)$ $=\lambda(x)\frac{\partial g}{\partial x_{i}}(x),\quad i=1,\dots,n,\quad\text{for all }x,$ (126)
then $f(x)$ is a transformation of $g(x)$; namely, there exists a function $T$ of one variable such that
$f(x)$ $=T(g(x))\quad\text{for all }x.$ (127)
Hence, in particular, the function $\lambda(x)$ satisfying (126) must be of the form:
$\lambda(x)$ $=\Lambda(g(x))\quad\text{for all }x,$ (128)
with some function $\Lambda$ of one variable.
If there is a function $T$ such that $f(x)=T(g(x))$, then by the chain rule, we have $\frac{\partial f}{\partial x_{i}}(x)=T^{\prime}(g(x))\frac{\partial g}{\partial x_{i}}(x)$ where $T^{\prime}$ is the derivative of $T$. Thus, by defining $\lambda(x)=T^{\prime}(g(x))$, we have $\frac{\partial f}{\partial x_{i}}(x)=\lambda(x)\frac{\partial g}{\partial x_{i}}(x)$. Intuitively, the lemma implies that the converse of the chain rule holds under additional conditions.